What is the demand for money when interest rates are zero and $Y=80?
b) If $Y= 80, what is the smallest value of money supply at which the interest rate is
zero?
c) Once the interest rate is zero, can the central bank continue to increase money
supply?
d) The United States experienced a long period of zero interest rates after 2009.
Can you find evidence in the text that the money supply continued to increase
over this period?
1
Expert's answer
2020-10-29T07:15:10-0400

(a) to determine Md (demand for money), we need the dependence between Md and i (interest rates)

For example if money demand is given by Md = $Y(0.25 - i)

Md=0.25*80=20 (i=0)

(b)the smallest value of money supply will be equal to Md (20, if we assume that Md = $Y(0.25 - i))

(c)yes, it can. For example by QE.

(d)the lower the interest rates, the greater the money supply. Evidence is on the picture